OpenStudy (anonymous):

Use the substitution method to solve the system of equations. Choose the correct ordered pair. 2x + 3y = 13 x = 2 A. (2, 4) B. (3, 2) C. (3, 3) D. (2, 3)

OpenStudy (anonymous):

2(2) + 3y = 13 4 + 3y = 13 3y = 9 y=3 x=2 (x,y) D. (2,3)